OpenStudy (anonymous):

6x^2-31x+5 divided (x-5)

OpenStudy (anonymous):

\[(6x^2-31x+5)\div(x-5)\]

OpenStudy (anonymous):

\[{6x^2-31+5 \over x-5}={(6x-1)(x-5) \over x-5}=6x-1\]

OpenStudy (anonymous):

\[6x ^{2}-31x+5 = 6x ^{2} -30x-x+5 =6x(x-5)-1(x-5) =(x-5)(6x-1)\] now, \[(x-5)(6x-1) /x-5 =6x-1\]